### v2

# Martin Boundary and the Nonlinear Multiplicative Stochastic Heat Equation in Weak Disorder

Abstract
We study the nonlinear multiplicative stochastic heat equation on Dirichlet spaces with white in time noise under weak disorder. We show that positive invariant random fields with uniformly bounded second moments are in one-to-one correspondence with bounded positive harmonic functions. The proof combines a remote past pullback construction with a uniqueness argument that applies a contraction inspired by chaos expansion. As a consequence, the space of invariant measures inherits geometric structure from the Martin Boundary. We further establish a small-noise Gaussian fluctuation result within each harmonic sector and show that the long-time behavior of solutions is completely determined by the Martin boundary data of the initial condition. These results reveal a direct connection between stochastic PDE dynamics and boundary theory in potential analysis.
1 Introduction
Let $(M,d,\mu)$ be a metric measure space equipped with a symmetric Dirichlet form $\mathcal{E}$. Denote the heat semigroup and $L^{2}(M)$-self-adjoint operator associated with $\mathcal{E}$ by $P_{t}$ and $\mathcal{L}$, respectively. We study the multiplicative stochastic heat equation, formally written as
$\partial_{t}u(t,x)=\mathcal{L}u(t,x)+\beta\,f(u(t,x))\,\dot{W}(t,x),\quad t>0,x\in M.$ (1)
Here, $\beta>0$ is the inverse temperature parameter, $f:\mathbb{R}\to\mathbb{R}$ is a Lipschitz function satisfying $u~||f(u)-f(\tilde{u})|\leq L_{f}$|u-$\tilde${u}| and $f(0)=0$, and $\dot{W}=\dot{W}(t,x)$ is a centered Gaussian noise with covariance kernel $R$ (the exact definition will be given in section 2). We will assume $R$ satisfies the following assumption for this paper.
Assumption 1.1.
Let $p_{t}$ denote the heat kernel, i.e. the kernel of the heat semigroup $P_{t}$, which we will assume exists. We say $R$ satisfies this assumption if there exists $\Lambda\in(0,\infty)$ such that
$\sup_{x,x^{\prime}\in M}\int_{0}^{\infty}\iint_{M^{2}}p_{t}(x,y)p_{t}(x^{\prime},y^{\prime})|R(y,y^{\prime})|dydy^{\prime}dt=\Lambda.$
We say any $f:M\to\mathbb{R}$ satisfying $\mathcal{L}f=0$ is a harmonic function.
Theorem 1.2.
Let $u_{0}\in L^{\infty}(M)$ and suppose $R$ satisfies Assumption 1.1. Let $u$ be the solution of (1) starting from $u_{0}$. The following statements hold.
  1. (i)
    If $(\beta L_{f})^{2}\Lambda<1$, then $u$ is in the weak disorder regime, that is,
    $\sup_{t\geq 0,x\in M}\mathbb{E}[u(t,x)^{2}]<\infty.$
  2. (ii)
    Let $(\beta L_{f})^{2}\Lambda<1$. There exists a one–to–one correspondence between positive invariant random fields $Z$ of (1) satisfying
    $\sup_{x\in M}\mathbb{E}[Z(x)^{2}]<+\infty$
    and bounded positive harmonic functions on $M$.
Remark 1.3 (Nontriviality of the assumption on $R$).
Assumption (1.1) has already been shown to be true in various settings, all non-compact. In Euclidean space $\mathbb{R}^{d}$ with $\mathcal{L}=\mathop{}\!\mathbin{\bigtriangleup}$ and $R$ bounded and compactly supported, it holds in $d\geq 3$. More interestingly, [BCO25, Lemma 8] showed that on Cartan-Hadamard manifolds with a negative section curvature upper bound and $\mathcal{L}=\mathop{}\!\mathbin{\bigtriangleup}_{M}$ the Laplace-Bemtrami operator, it holds for $R$ with Riesz-type covariance. For the special case $M=\mathbb{H}^{d}$, [GO25, Lemma 2.4] identified the sufficient condition $|R(x,y)|\leq Cd(x,y)^{-\alpha},$ where $\alpha>1$ for the assumption to hold. These last two examples have very interesting Martin boundaries at infinity, and the verification comes down to exponential in time heat kernel bounds on diagonal. This means that extensions to non-manifold settings is also simply a matter of constructing appropriate $R$ and looking at heat kernel behavior. Notably, on regular metric trees, it should be easy to verify that the assumption holds for space-time white noise ($R(x,y)=\delta_{x=y}$) using known heat kernel upper bounds[Cro08, ALW11, Hon24]. This is expected to generalize to all non-amenable metric graphs. While these are all examples where $\mathcal{E}$ is local, the theorem does not require that, and indeed we expect that there are examples of nonlocal spaces and appropriate $R$ satisfying the fundamental assumption.
Remark 1.4 (On the moment assumption).
The uniform pointwise second-moment bound imposed in Theorem 1.2 is the natural condition corresponding to the weak disorder/diffusive regime. In the linear case $f(u)=u$, it indicates that the partition function of the associated polymer has uniformly bounded second moments, reflecting the absence of localization and the persistence of diffusive behavior. The class of invariant random fields considered here therefore corresponds to physically relevant stationary polymer states in the weak disorder phase.
Next is a fluctuation result which differs across harmonic sectors.
Theorem 1.5 (First order Gaussian fluctuation in $\beta\to 0$ limit of invariant fields).
For $\beta$ such that $(\beta L_{f})^{2}\Lambda<1$ and $h\in\mathrm{Har}^{+}_{b}(M)$, let $Z^{h}_{\beta}$ be the time-invariant random field of (1) associated with $h$ at inverse temperature $\beta$. Define a centered Gaussian field
$G_{h}(x):=\int_{-\infty}^{0}\int_{M}p_{-s}(x,y)f(h(y))W(dy,ds).$
We have
$\frac{Z^{h}_{\beta}-h}{\beta}\stackrel{{\scriptstyle\beta\downarrow 0}}{{\longrightarrow}}G_{h}\quad\text{ in }L^{2}(\Omega)\text{ as random fields}.$
Finally, we show that the invariant measures identified in Theorem 1.2 completely determine the long time dynamics of the solution $u$ under mild assumptions on the initial condition.
Theorem 1.6.
Suppose $u_{0}\in L^{\infty}(M)$ satisfies $\lim_{t\to\infty}P_{t}u_{0}=h\in\mathrm{Har}^{+}_{b}(M)$ in $C(M)$, $(\beta L_{f})^{2}\Lambda<1$. Let $Z^{h}_{\beta}$ denote the unique invariant field of (1) associated with $h$. Then $u$ solving (1) starting from $u_{0}$ converges in $L^{2}(\Omega)$ to $Z^{h}_{\beta}$.
The paper is organized as follows. We will discuss existing literature to finish off the introduction. In section 2, we list some useful preliminary definitions and fix some notation. We will then move on to prove the main theorem in section 3. Item (i) requires almost no new mathematics, so most of the section is reserved for proving item (ii), the classification result. For that, we first show existence given a bounded positive harmonic $h$ result using an argument originating in [GL20] (see also [GHL25, Che+25]), then we conclude by showing uniqueness per $h$ by a chaos expansion inspired contraction. We then move onto proving Theorems 1.5 and 1.6 in sections 4 and 5, respectively. In the appendix, we discuss the implications of our results by it’s connection to Martin boundary theory.
1.1 Literature Review
Equation (1) is interesting on its own[Mol91] and for it’s connection to various other areas of probability theory and mathematical physics. In the linear case $f(u)=u$, it is the partition function of directed polymers in a random environment[ACQ10, CY04, CSZ20, DZ22, DGL23] and gives rise to the Cole–Hopf solution of the KPZ equation [Kar87, KPZ86]. It also has direct connections to the stochastic Burgers equation [BCJ94, BG97, CM94] and to Majda’s model of shear-layer flow in turbulent diffusion [Maj93]. In the last few decades, there has been significant progress in understanding fundamental properties of (1) on flat spaces $\mathbb{R}^{d}$ and $\mathbb{T}^{d}$, including intermittency (see, e.g., [CK19, Kho14] and references therein), as well as energy landscapes and fluctuations [DGK23, GHL25, GK24, GL20]. In particular, invariant measures have been heavily studied[Che+25, Dun+18, DM24, DS24, GQ24, Gro+23] due to their connections with fluctuation theory and other deeper properties. Notably, it has been shown that nontrivial invariant fields exist in the weak disorder/diffusive regime and this has been connected to some of the aforementioned fluctuation results.
In contrast, the literature on non-Euclidean settings has only begun to develop recently, almost all of it for the linear case and focused on well-posedness and/or pointwise moment or almost sure asymptotics. On compact Riemannian manifolds, [TV02, CO25] showed that strong disorder/localization always occurs, the latter also dealing with difficulties introduced by geometry for well-posedness if one allows measure-valued initial conditions. Well-posedness theory for bounded initial conditions, on the other hand, have been extended to stranger geometries such as sub-Riemannian manifolds[Bau+23] and fractals [Bau+25, HY16]. All the above works have also dealt with moment estimates, which are studied for their connection to intermittency/localization. Most recently, the effect of negative curvature on moment and pointwise asymptotics have begun to be explored. In [GX25, GWX25], it was shown that for the linear case with noise independent of time, moment asymptotics in Hyperbolic space do not behave differently compared to the Euclidean case, but almost sure pointwise asymptotics grow at a different rate. For the white in time noises (the setting of this paper), [BCO25] showed that for a natural class of noises, negative curvature generates a weak disorder regime. This result was sharpened in [GO25], who showed that all noises with $R$ polynomially decaying at rate $r>1$ gives rise to such a regime. Remarkably, they discovered that if $0<r<1$ instead, in place of a weak disorder regime a new regime where strong disorder ($\lim_{t\to+\infty}\mathbb{E}[u(t,x)^{2}]=\infty$) still holds, but the rate of growth is sub-exponential as opposed to exponential. This phenomenon has never been observed before in any models of the linear equation with white in time noise.
Motivated by these developments, one may ask whether the invariant measures of (1) in the weak disorder (diffusive) regime is connected to the structure of harmonic functions on the underlying space. In Euclidean settings, this question is trivial due to Liouville’s theorem, which forces all bounded harmonic functions to be constant and thus yields a unique spatially homogeneous stationary state. In contrast, on non-compact spaces with nontrivial Martin boundary, such as negatively curved manifolds[Anc87, AS85] or trees[Cro08], the abundance of bounded positive harmonic functions suggests the possibility of a rich structure of invariant measures reflecting boundary geometry. Theorem 1.2 answers this question in a precise and complete form: invariant random fields with uniformly bounded second moments are in one-to-one correspondence with bounded positive harmonic functions. In other words, the invariant measures inherit the structure possessed by the bounded positive harmonic functions, establishing a connection with Martin boundary theory. To our knowledge, this provides the first example of a classification of invariant measures for a stochastic PDE in terms of the harmonic (or boundary) structure of the underlying space.
1.2 Acknowledgements
The author thanks Le Chen and Alexander Dunlap for helpful discussions which motivated this work, as well as Fabrice Baudoin for pointing out an error in a previous version. This work was supported by a research grant (VIL73729) from Villum Fonden.
2 Preliminaries
We will first fix some notation and conventions used for the rest of the paper.
  • We follow the usual convention and use $C,c$. to denote generic constants that are independent of quantities of interest. The exact values of these constants may change from line to line.
  • For $t>0,x,y\in M$, $P_{t}$ will denote the heat semigroup and $p_{t}(x,y)$ will denote the heat kernel on $M$.
  • $\partial_{M}$ will denote the Martin boundary of $M$.
  • $\mathrm{Har}^{+}_{b}(M)$ will denote the convex cone of bounded harmonic functions on $M$.
  • dzdz will denote integration with respect to the $\mu$ on $M$ in the variable $z$.
We now introduce centered Gaussian noises with covariance kernel $R$ and rigorously define the notion of a solution. Any positive-definite function $R$ on $M$ defines an inner product by
$\braket{\phi,\psi}_{R}:=\int_{0}^{\infty}\iint_{M^{2}}\phi(t,x)R(x,y)\psi(t,y)dxdydt.$
We denote the Hilbert space of space-time functions corresponding to $\braket{\cdot,\cdot}_{R}$ by $\mathcal{H}_{R}$. On a probability space $(\Omega,\mathcal{G},\P)$, the noise $\dot{W}$ is defined as the isonormal Gaussian process[Nua95] over $\mathcal{H}_{R}$ such that $\mathcal{G}=\sigma(\dot{W})$. We will not denote its $R$-dependence as there is no concern for confusion.
The stochastic integral of a random field $F$ with respect to $W$, denoted by
$I(F):=\int_{0}^{\infty}\int_{M}F(t,x)W(dx,dt),$
is defined for any $F$ such that
$\mathbb{E}[I(F)^{2}]=\int_{0}^{\infty}\iint_{M^{2}}\mathbb{E}[F(t,x)F(t,y)]R(x,y)dxdydt<\infty.$
We call such $F$ the $W$-integrable random fields. For any $F,\widetilde{F}$ which are $W$-integrable, we have
$\mathbb{E}[I(F)I(\widetilde{F})]=\int_{0}^{\infty}\iint_{M^{2}}\mathbb{E}[F(t,x)\widetilde{F}(t,y)]R(x,y)dxdydt.$
We call the above formulas Itô–Walsh isometry.
Definition 2.1.
A process $u=\{u(t,x);(t,x)\in\mathbb{R}_{+}\times M\}$ is called a random field solution of (1) in the Itô-Walsh sense [Wal86] if the following conditions are met:
  1. 1.
    $u$ is adapted;
  2. 2.
    $u$ is jointly measurable with respect to $\mathcal{B}(\mathbb{R}_{+}\times M)\otimes\mathcal{G}$;
  3. 3.
    $\mathbf{E}(u(t,x)^{2})<\infty$ for all $(t,x)\in\mathbb{R}_{+}\times M$;
  4. 4.
    The function $(t,x)\to u(t,x)$ is continuous in $L^{2}(\Omega)$;
  5. 5.
    $u$ satisfies
    $u(t,x)=P_{t}u_{0}(x)+\beta\int_{0}^{t}\int_{M}p_{t-s}(x,y)f(u(s,y))W(ds,dy)$
    almost surely for all $(t,x)\in\mathbb{R}_{+}\times M$.
Remark 2.2.
We note that by Itô–Walsh isometry, the second moment of the solution $u$ starting from $u_{0}$ has the useful representation
$\mathbb{E}[u(t,x)]=P_{t}u_{0}(x)^{2}+\beta^{2}\int_{0}^{t}\iint_{M^{2}}p_{t-s}(x,y)p_{t-s}(x,y^{\prime})R(y,y^{\prime})\mathbb{E}[f(u(s,y))f(u(s,y^{\prime}))]dydy^{\prime}ds.$
3 Proof of Main Theorem
The proof of item (i) follows from repeatedly applying Itô-Walsh isometry into the mild formulation (5) along with $\mathbb{E}[f(u(s,y))f(u(s,y^{\prime}))]\leq L_{f}^{2}\sup_{y\in M}\mathbb{E}[u(s,y)^{2}]$ to obtain a Neumann-type series (see [CD15, CK19] for this iteration being used to obtain well-posedness), and allowing time to go to infinity yields
$\mathbb{E}[u(t,x)^{2}]\leq\norm{u_{0}}_{\infty}\sum_{k\geq 0}((\beta L_{f})^{2}\Lambda)^{k},$
and the RHS geometric series converges iff $(\beta L_{f})^{2}\Lambda<1$.
For item (ii), we first show that the mean of an invariant field must be harmonic. It is simple: taking expectation of (5) gives us
$\mathbb{E}[u(t,x)]=\int_{M}p_{t}(x,y)u_{0}(y)dy,$
i.e. the expectation of (1) started from $u_{0}$ is just the solution of the homogeneous heat equation starting from $u_{0}$. If a field is invariant, then its first moment must also be invariant, and the only functions invariant under the heat semigroup are precisely the harmonic functions. This means that any invariant measure must be associated to some harmonic function $h$. Also, this immediately rules out the possibility that two harmonic functions are associated with the same invariant measure, as their means, the harmonic functions, are not the same.
We are thus left to prove two steps to finish the proof of item (ii) and therefore the entire theorem: an existence step, for which we adapt the method of [GL20] to solutions of (1) starting from bounded harmonic functions, and a uniqueness step, for which we will again utilize chaos expansion.
3.1 Existence by Backward Time Shifts
This version of the argument first showcased in [GL20] is inspired by the version [GHL25, Theorem 2.10].
Theorem 3.1 (Pullback construction from harmonic initial condition).
Suppose $\beta<\Lambda^{-\frac{1}{2}}$. For $h\in\mathrm{Har}^{+}_{b}(M)$ and $K>0$, we define $u^{h,K}(t,x)$ as the solution $K$-backward shifted equation starting from $h$, that is,
$u^{h,K}(t,x)=h(x)+\beta\int_{-K}^{t}\int_{M}p_{t}(x,y)f(u^{h,K}(s,y))W(ds,dy).$
Then:
  1. (i)
    For every fixed $t\in\mathbb{R}$ and $x\in M$, the family $\{u^{K,h}(t,x)\}_{K>-t}$ is Cauchy in $L^{2}(\Omega)$. Hence there exists a limit
    $\widetilde{Z}^{h}(t,x):=\lim_{K\to\infty}u^{K,h}(t,x)\quad\text{in }L^{2}(\Omega).$
  2. (ii)
    The random field $\widetilde{Z}^{h}(t,x)$ is stationary in time. In particular, the time-zero marginal
    $Z^{h}(x):=\widetilde{Z}^{h}(0,x)$
    is invariant in law under the Markov evolution of the stochastic heat equation (1).
  3. (iii)
    For all $x\in M$,
    $\mathbb{E}[Z^{h}(x)]=h(x).$
Proof.
For item (i), we first fix $t\in\mathbb{R}$ and let $K^{\prime}>K>-t$, and for them define the difference
$\delta^{K,K^{\prime}}(t,x):=u^{K^{\prime},h}(t,x)-u^{K,h}(t,x).$
We also define
$A(K,K^{\prime}):=\sup_{-K^{\prime}<-K<t,x\in M}\mathbb{E}[\delta^{K,K^{\prime}}(t,x)^{2}].$
By the mild formulation (5), we have
$\delta^{K,K^{\prime}}(t,x)=$ $\beta I(-K,t)+\beta I(-K^{\prime},-K),\text{ where }$
$I(-K,t):=$ $\int_{-K}^{t}\int_{M}p_{t-s}(s,y)[f(u^{K^{\prime},h}(s,y))-f(u^{K,h}(s,y))]W(ds,dy),$
$I(-K^{\prime},-K):=$ $\int_{-K^{\prime}}^{-K}\int_{M}p_{t-s}(s,y)f(u^{K^{\prime},h}(s,y))W(ds,dy).$
We will now bound the second moments of these stochastic integrals. By Itô–Walsh isometry,
$\mathbb{E}[I(-K,t)^{2}]=$ $\int_{-K}^{t}\iint_{M^{2}}p_{t-s}(x,y)p_{t-s}(x,y^{\prime})R(y,y^{\prime})$
$\times\mathbb{E}[\{f(u^{-K^{\prime}}(s,y))-f(u^{-K}(s,y))\}\{f(u^{-K^{\prime}}(s,y^{\prime}))-f(u^{-K}(s,y^{\prime}))\}]dydy^{\prime}ds$
$\leq$ $\int_{-K}^{t}\iint_{M^{2}}p_{t-s}(x,y)p_{t-s}(x,y^{\prime})R(y,y^{\prime})dydy^{\prime}ds$
$\times\sup_{y\in M}\mathbb{E}[\{f(u^{-K^{\prime}}(s,y))-f(u^{-K}(s,y))\}^{2}]$
$\leq$ $\Lambda L_{f}^{2}A(K,K^{\prime}).$
In the last step, we used Assumption 1.1 and $g$ being Lipschitz. Similarly,
$\mathbb{E}[I(-K^{\prime},-K)^{2}]=$ $\int_{-K^{\prime}}^{-K}\iint_{M^{2}}p_{t-s}(x,y)p_{t-s}(x,y^{\prime})R(y,y^{\prime})\mathbb{E}[f(u^{K^{\prime}}(s,y))f(u^{K^{\prime}}(s,y^{\prime}))]dydy^{\prime}ds$
$\leq$ $L_{f}^{2}\sup_{s>-K^{\prime},y\in M}\mathbb{E}[u^{-K^{\prime}}(s,y)^{2}]\int_{-K^{\prime}}^{-K}\iint_{M^{2}}p_{t-s}(x,y)p_{t-s}(x,y^{\prime})R(y,y^{\prime})dydy^{\prime}ds$
$=$ $L_{f}^{2}\sup_{s>-K^{\prime},y\in M}\mathbb{E}[u^{-K^{\prime}}(s,y)^{2}]\left[\int_{-\infty}^{-K}-\int_{-\infty}^{-K^{\prime}}\right]\iint_{M^{2}}p_{t-s}(x,y)p_{t-s}(x,y^{\prime})R(y,y^{\prime})dydy^{\prime}ds$
$=$ $L_{f}^{2}\sup_{s>-K^{\prime},y\in M}\mathbb{E}[u^{-K^{\prime}}(s,y)^{2}](F(K)-F(K^{\prime})).$
By Assumption 1.1, the function
$F(K)=\int_{-\infty}^{-K}\iint_{M^{2}}p_{t-s}(x,y)p_{t-s}(x,y^{\prime})R(y,y^{\prime})dydy^{\prime}ds$
is Cauchy in $K$ as $K\uparrow\infty$ for K<t-K<t. Combining, we have
$\mathbb{E}[\delta^{K,K^{\prime}}(t,x)^{2}]\leq(\beta L_{f})^{2}\Lambda A(K,K^{\prime})+(\beta L_{f})^{2}\sup_{s>-K^{\prime},y\in M}\mathbb{E}[u^{-K}(s,y)^{2}](F(K)-F(K^{\prime})).$
Taking sup over $t,x$ in LHS and noting that $\sup_{s>-K^{\prime},y\in M}\mathbb{E}[u^{-K}(s,y)^{2}]\leq C$ by Assumption 1.1 gives
$(1-(\beta L_{f})^{2}\Lambda)A(K,K^{\prime})\leq(\beta L_{f})^{2}C(F(K)-F(K^{\prime})\iff\frac{(\beta L_{f})^{2}C}{1-(\beta L_{f})^{2}\Lambda}[F(K)-F(K^{\prime})].$
Since $F(K)$ is Cauchy in large $K$, this finishes the proof of item (i).
Now we move to item (ii). Fix $\tau>0$ and $K>\tau$. By the flow property,
$u^{K,h}(\tau,\cdot)=S_{\tau}\big(u^{K,h}(0,\cdot)\big),$
where $S_{\tau}$ denotes the semigroup corresponding to (1) acting on probability measures over $C_{\mathrm{loc}}(M)$.
Since the noise is stationary under time shifts, the law of $u^{K,h}(\tau,\cdot)$ coincides with that of $u^{K-\tau,h}(0,\cdot)$. Therefore
$S_{\tau}\big(u^{K,h}(0,\cdot)\big)\stackrel{{\scriptstyle\mathrm{law}}}{{=}}u^{K-\tau,h}(0,\cdot).$
Letting $K\to\infty$ and using the $L^{2}$ convergence established above, we obtain
$S_{\tau}Z^{h}\stackrel{{\scriptstyle\mathrm{law}}}{{=}}Z^{h}.$
Hence $Z^{h}$ is invariant in law. The same argument shows that the full field $\widetilde{Z}^{h}(t,\cdot)$ is stationary in time.
Finally, we show item (iii). Taking expectations in the mild formulation gives
$\mathbb{E}[u^{K,h}(t,x)]=P_{t+K}h(x)=h(x).$
Passing to the limit yields item (iii). This completes the proof. ∎
3.2 Uniqueness by Chaos Expansion
In Theorem 3.1, we obtained for every $h\in\mathrm{Har}^{+}_{b}(M)$ an invariant random field of the PAM dynamic which has mean $h$. We will now show this random field is unique for each choice of such $h$.
Theorem 3.2 (Uniqueness of stationary solutions in the weak-disorder class).
Assume $(\beta L_{f})^{2}\Lambda<1$ and let $V^{1},V^{2}$ be two stationary mild solutions of (1) with the same first moment such that
$\sup_{t\in\mathbb{R}}\sup_{x\in M}\mathbb{E}[|V^{i}(t,x)|^{2}]<\infty,\quad i\in\set{1,2}$.
Then $V^{1}\equiv V^{2}$ almost surely.
Proof.
Define $\delta(t,x):=V^{1}(t,x)-V^{2}(t,x)$. Since neither $V^{1}$ nor $V^{2}$ have their laws change in time, $\delta$ is another stationary solution and since $V^{1}$ and $V^{2}$ have the same pointwise mean, $\mathbb{E}[\delta(t,\cdot)]\equiv 0$.
The theorem is proven if $\delta\equiv 0$. Since $\delta$ is time-stationary, we extend it to all $t\in\mathbb{R}$. Thus the back-shifted mild formulation of $\delta$ at 0 from T-T yields
$\delta(0,x)=[P_{T}\delta(-T,\cdot)](x)+\beta\int_{-T}^{0}\int_{M}p_{-s}(x,y)[f(V^{1}(y))-f(V^{2}(y))]W(ds,dy)$
We will first show
$A_{T}(x):=P_{T}\delta(-T,\cdot),\qquad\lim_{T\uparrow\infty}A_{T}(x)=0.$
Define the remote past filtration $\mathcal{F}_{-T}:=\sigma(W((-\infty,-T]\times M))$. Since the Walsh integral over (T,0](-T,0] has expectation 0 when conditioned on $\mathcal{F}_{-T}$, we have
$\mathbb{E}[\delta(0,\cdot)|\mathcal{F}_{-T}]=P_{T}\delta(-T,\cdot)](x).$
Sending $T\uparrow\infty$ in the above and applying Kolmogorov’s 0–1 law ($\mathcal{F}_{-\infty}:=\cap_{T\geq 0}\mathcal{F}_{-T}=\emptyset$) along with the martingale convergence theorem backwards gives us (3.2).
Now let $A(x,y):=\mathbb{E}[\delta(0,x)\delta(0,y)]$. If $A\equiv 0$, then $\mathbb{E}[(\delta(0,x)-\delta(0,y))^{2}]=0$ for all $x,y$, hence $\delta\equiv 0$ almost surely. The rest of this proof is thus dedicated to showing $A\equiv 0$.
By Itô-Walsh isometry, we have
$A(x,x^{\prime})=$ $\mathbb{E}[A_{T}(x)A_{T}(x^{\prime})]$
$+\beta^{2}\int_{-T}^{0}\iint_{M^{2}}p_{-s}(x,y)p_{-s}(x,y^{\prime})R(y,y^{\prime})\mathbb{E}[\{f(V^{1}(y))-f(V^{2}(y))\}\{f(V^{1}(y^{\prime}))-f(V^{2}(y^{\prime}))\}]dydy^{\prime}ds.$
$\leq$ $E[A_{T}(x)A_{T}(x^{\prime})]$
$+\beta^{2}\int_{-T}^{0}\iint_{M^{2}}p_{-s}(x,y)p_{-s}(x,y^{\prime})R(y,y^{\prime})\sup_{z\in M}\mathbb{E}[|f(V^{1}(z))-f(V^{2}(z))|^{2}]dydy^{\prime}ds$
$\leq$ $E[A_{T}(x)A_{T}(x^{\prime})]$
$+\beta^{2}\int_{-T}^{0}\iint_{M^{2}}p_{-s}(x,y)p_{-s}(x,y^{\prime})R(y,y^{\prime})L_{f}^{2}\norm{A}_{\infty}dydy^{\prime}ds$
$\leq$ $E[A_{T}(x)A_{T}(x^{\prime})]+(\beta L_{f})^{2}\Lambda\norm{A}_{\infty},$
where $\norm{A}_{\infty}:=\sup_{x,y\in M}|A(x,y)|$. Taking $\sup_{x,y\in M}$ in LHS and sending $T\uparrow\infty$, we have by (3.2)
$\norm{A}_{\infty}\leq(\beta L_{f})^{2}\Lambda\norm{A}_{\infty},$
which forces $\norm{A}_{\infty}=0$ if $(\beta L_{f})^{2}\Lambda<1$. This concludes the proof. ∎
4 Proof of Theorem 1.5
By Itô-Walsh Isometry, we have
$\mathbb{E}[G_{h}(x)G_{h}(x^{\prime})]=\int_{0}^{\infty}\iint_{M^{2}}p_{t}(x,y)p_{t}(x^{\prime},y^{\prime})R(y,y^{\prime})f(h(y))f(h(y^{\prime}))dydy^{\prime}t.$
Define two quantities
$\Delta_{\beta}(x):=Z^{h}_{\beta}(x)-h(x),\qquad M_{\beta}:=\sup_{x\in M}\mathbb{E}[\Delta_{\beta}(x)^{2}].$
Let $Z^{h}_{\beta}(s,\cdot)$ be $Z^{h}_{\beta}$ realized as a space-time and time-stationary field. We will show that
  1. (i)
    $M_{\beta}=O(\beta^{2})$ as $\beta\downarrow 0$,
  2. (ii)
    $\mathbb{E}\big[\big|\frac{\Delta_{\beta}(x)}{\beta}-G_{h}(x)\big|^{2}\big]\leq CM_{\beta}$.
Proving both of the above would conclude the proof.
By Itô-Walsh isometry,
$\mathbb{E}[\Delta_{\beta}(x)^{2}]=$ $\beta^{2}\int_{-\infty}^{0}\iint_{M^{2}}p_{-s}(x,y)p_{-s}(x,y^{\prime})R(y,y^{\prime})\mathbb{E}[f(Z^{h}_{\beta}(s,y))f(Z^{h}_{\beta}(s,y^{\prime}))]dydy^{\prime}dt,$
$\mathbb{E}[f(Z^{h}_{\beta}(s,y))f(Z^{h}_{\beta}(s,y^{\prime}))]\leq$ $L_{f}^{2}\sup_{x\in M}\mathbb{E}[Z^{h}_{\beta}(x)^{2}]=L_{f}^{2}\mathbb{E}[(\Delta_{\beta}(x)+h(x))^{2}]=L_{f}^{2}(M_{\beta}+\norm{h}^{2}_{\infty}),\text{ thus }$
$\mathbb{E}[\Delta_{\beta}(x)^{2}]\leq$ $\beta^{2}L_{f}^{2}\Lambda M_{\beta}+\beta^{2}L_{f}^{2}\Lambda\norm{h}^{2}_{\infty}.$
Taking sup in $x$ in LHS gives us
$M_{\beta}\leq\frac{(\beta L_{f})^{2}\Lambda\norm{h}^{2}_{\infty}}{1-(\beta L_{f})^{2}\Lambda}=O(\beta^{2}),\quad\beta\downarrow 0.$
This proves item (i). For item (ii), we have simply
$\mathbb{E}\left[\left(\frac{\Delta_{\beta}}{\beta}-G_{h}(x)\right)^{2}\right]$
$=$ $\int_{0}^{\infty}\iint_{M^{2}}p_{t}(x,y)p_{t}(x,y^{\prime})R(y,y^{\prime})\mathbb{E}[(f(Z^{h}_{\beta}(s,y)-f(h(y)))(f(Z^{h}_{\beta}(s,y^{\prime})-f(h(y^{\prime})))]dydydt$
$\leq$ $\Lambda\sup_{z\in M}\mathbb{E}[(f(Z^{h}_{\beta}(s,z)-f(h(z)))^{2}]\leq L_{f}^{2}\Lambda M_{\beta}.$
Remark 4.1.
Even at first order in $\beta$, the limiting covariance (4) depends on the sector $h$ through the profile $f(h)$; in particular, distinct $h$ typically yield distinct Gaussian limits.
5 Proof of Theorem 1.6
We realize $Z^{h}$ as a time-stationary field and set $\delta(t,x):=u(t,x)-Z^{h}(t,x)$. By (5),
$\delta(t,x)=P_{t}(u_{0}-h)(x)+\beta\int_{0}^{t}\int_{M}p_{t-s}(x,y)\big(f(u(s,y))-f(Z^{h}(s,y))\big)\,W(ds,dy).$
Applying Itô isometry and using $f$ is Lipschitz, we obtain
$\mathbb{E}[\delta(t,x)^{2}]\leq$ $|P_{t}(u_{0}-h)(x)|^{2}$
$+(\beta L_{f})^{2}\int_{0}^{t}\iint_{M^{2}}p_{t-s}(x,y)p_{t-s}(x,y^{\prime})|R(y,y^{\prime})|\mathbb{E}[|\delta(s,y)\delta(s,y^{\prime})|]\,dydy^{\prime}\,ds.$
Using $|ab|\leq\frac{1}{2}(a^{2}+b^{2})$ and taking supremum over $x\in M$, we obtain
$M(t)\leq a(t)+(\beta L_{f})^{2}\Lambda\sup_{s\leq t}M(s),$
where
$M(t):=\sup_{x\in M}\mathbb{E}[\delta(t,x)^{2}],\qquad a(t):=\sup_{x\in M}|P_{t}(u_{0}-h)(x)|^{2}.$
Since $a(t)\to 0$ by assumption and $(\beta L_{f})^{2}\Lambda<1$, we conclude
$\sup_{s\leq t}M(s)\leq\frac{1}{1-(\beta L_{f})^{2}\Lambda}\sup_{s\leq t}a(s),$
and hence $M(t)\to 0$ as $t\to\infty$.
Remark 5.1.
The assumption that $P_{t}u_{0}$ converges uniformly to a harmonic function $h$ is natural in geometrically non-compact settings. In particular, on spaces with a well-understood Martin boundary (e.g. hyperbolic spaces and more generally negatively curved manifolds), bounded functions which admit a continuous extension to the boundary have harmonic limits given by the corresponding Poisson/Martin representation. In such cases, the above proposition shows that the long-time behavior of the stochastic equation is completely determined by the boundary data of the initial condition.
Appendix A Appendix: Connections to Martin Boundary
Martin boundary and boundary integral representation
For a detailed discussion of Martin boundary theory, see [FOT11].
Assume that the Markov process associated with $\mathcal{E}$ is transient and admits a Green kernel
$G(x,y):=\int_{0}^{\infty}p_{t}(x,y)\,dt<\infty\quad\text{for $\mu$-a.e.\ $x\neq y$.}$
Fix a base point $o\in M$. The Martin kernel is defined by
$K(x,y):=\frac{G(x,y)}{G(o,y)}.$
The Martin compactification $\overline{M}^{\,M}$ is the smallest compactification for which the maps $y\mapsto K(\cdot,y)$ extend continuously, and the Martin boundary is $\partial_{M}M:=\overline{M}^{\,M}\setminus M$.
Under standard hypotheses (e.g. those ensuring solvability of the Dirichlet problem at infinity and the boundary Harnack principle in concrete models), every $h\in\mathrm{Har}_{b}^{+}(M)$ admits a boundary integral representation
$h(x)=\int_{\partial_{M}M}K(x,\xi)\,\nu(d\xi),$ (2)
for a finite positive measure $\nu$ on $\partial_{M}M$; moreover extremal rays of $\mathrm{Har}_{b}^{+}(M)$ correspond to minimal Martin boundary points.
Remark A.1.
In classical negatively curved manifolds one often identifies $\partial_{M}M$ with the geometric boundary, while for trees (and more generally many transient graphs) $\partial_{M}M$ identifies with the space of ends. We keep the Martin boundary formulation throughout since it is intrinsic to the Markov semigroup and also accommodates non-local (jump) generators.
Equivariance under automorphisms of the measure, heat kernel, and noise
Let $\mathcal{G}$ be a group acting measurably on $M$. We assume the action preserves the reference measure $\mu$ and intertwines the semigroup:
$p_{t}(gx,gy)=p_{t}(x,y)\quad\text{for all }g\in\mathcal{G},\ t>0,$
and that the spatial covariance kernel $R$ of the driving noise is $\mathcal{G}$-invariant:
$R(gx,gy)=R(x,y)\quad\text{for all }g\in\mathcal{G}.$
If $g\in\mathcal{G}$ and $h$ is harmonic, define
$(g\cdot h)(x):=h(g^{-1}x).$
Then $g\cdot h$ is harmonic, and harmonic measures transform according to
$\omega_{gx}=g_{\#}\omega_{x},$
which implies
$K(gx,\xi)=K(x,g^{-1}\xi).$
Thus the action of isometries corresponds to pushforward of boundary data. We now show (1) is equivariant under the action of $\mathcal{G}$ by generalizing the argument of [Che+21, Lemma 7.1].
Proposition A.2 (Isometry equivariance in mild form).
Define the $\mathcal{G}$-action on functions (and random fields) by
$(g\cdot\varphi)(x):=\varphi(g^{-1}x),\qquad g\in\mathcal{G}.$
Assume the spatial covariance kernel $f$ of the noise is $G$-invariant, i.e.
$R(gx,gy)=R(x,y)\qquad\text{for all }g\in G,\ x,y\in\mathbb{H}^{d}.$
Define the transformed noise by
$W^{g}(\phi):=W(\phi^{g}),\qquad\phi^{g}(t,x):=\phi(t,gx),$
so that formally $\dot{W}^{g}(t,x)=\dot{W}(t,g^{-1}x)$.
Let $u$ solve (1) starting from $u_{0}\in L^{\infty}(M)$. Then for each $g\in\mathcal{G}$ the shifted field
$u^{g}(t,x):=(g\cdot u(t,\cdot))(x)=u(t,g^{-1}x)$
is the mild solution with initial condition $g\cdot u_{0}$ driven by noise $W^{g}$. In particular, since $W^{g}\stackrel{{\scriptstyle\mathrm{law}}}{{=}}W$ by $\mathcal{G}$-invariance of $R$, the law of $u^{g}$ coincides with the law of the solution driven by $W$ started from $g\cdot u_{0}$.
Proof.
We first show that the assumption on $R$ gives $\mathcal{G}$-invariance of noise. Indeed, for test functions $\phi,\psi\in C_{c}^{\infty}(\mathbb{R}_{+}\times M)$,
$\mathbb{E}[W^{g}(\phi)W^{g}(\psi)]=\mathbb{E}[W(\phi^{g})W(\psi^{g})]=\int\phi(t,gx)\psi(t,gy)\,R(x,y)\,dx\,dy\,dt=\int\phi(t,x)\psi(t,y)\,R(x,y)\,dx\,dy\,dt,$
where we used the change of variables $x\mapsto g^{-1}x$, $y\mapsto g^{-1}y$ and invariance of $R$ and dxdydxdy. Hence $W^{g}\stackrel{{\scriptstyle\mathrm{law}}}{{=}}W$.
Fix $g\in\mathcal{G}$ and set $u^{g}(t,x):=u(t,g^{-1}x)$. Using invariance of the heat kernel under isometries,
$p_{t}(gx,gy)=p_{t}(x,y),\qquad t>0,\ x,y\in\mathbb{H}^{d},$ (3)
and $\mathcal{G}$-invariance of $\mu$, we obtain
$(P_{t}u_{0})(g^{-1}x)=\int_{M}p_{t}(g^{-1}x,y)\,u_{0}(y)\,dy=\int_{M}p_{t}(x,gy^{\prime})\,u_{0}(gy^{\prime})\,dy^{\prime}=(P_{t}(g\cdot u_{0}))(x),$
where we changed variables $y=gy^{\prime}$ and used (3).
For the stochastic integral term, define the integrand
$\Phi_{t,x}(s,y):=\mathbf{1}_{[0,t]}(s)\,p_{t-s}(g^{-1}x,y)\,u(s,y),$
so that
$\int_{0}^{t}\int_{M}p_{t-s}(g^{-1}x,y)\,u(s,y)\,W(ds,dy)=W(\Phi_{t,x}).$
Now define
$\widetilde{\Phi}_{t,x}(s,y^{\prime}):=\mathbf{1}_{[0,t]}(s)\,p_{t-s}(x,y^{\prime})\,u^{g}(s,y^{\prime})=\mathbf{1}_{[0,t]}(s)\,p_{t-s}(x,y^{\prime})\,u(s,g^{-1}y^{\prime}).$
Using $p_{t-s}(g^{-1}x,y)=p_{t-s}(x,gy)$ from (3), we have
$\Phi_{t,x}(s,y)=\widetilde{\Phi}_{t,x}(s,gy),$
i.e. $\Phi_{t,x}=(\widetilde{\Phi}_{t,x})^{g}$ with the convention $\psi^{g}(s,y):=\psi(s,gy)$. Therefore, by definition of the transformed noise $W^{g}$,
$W(\Phi_{t,x})=W\big((\widetilde{\Phi}_{t,x})^{g}\big)=W^{g}(\widetilde{\Phi}_{t,x})=\int_{0}^{t}\int_{M}p_{t-s}(x,y^{\prime})\,f(u^{g}(s,y^{\prime}))\,W^{g}(ds,dy^{\prime}).$
Combining the transformed deterministic and stochastic terms yields
$u^{g}(t,x)=(P_{t}(g\cdot u_{0}))(x)+\beta\int_{0}^{t}\int_{M}p_{t-s}(x,y)\,f(u^{g}(s,y))\,W^{g}(ds,dy),$
which is exactly (5) with initial condition $g\cdot u_{0}$ driven by $W^{g}$.∎
Corollary A.3 (Equivariance of invariant fields and invariant measures).
Assume the hypotheses of Proposition A.2. Let $\mathcal{S}_{t}^{W}$ denote the Markov solution map at time $t$ for (1) driven by noise $W$, i.e. $\mathcal{S}_{t}^{W}(u_{0})$ is the (mild) solution at time $t$ started from $u_{0}$.
  1. (i)
    (Equivariance of the dynamics.) For every $g\in\mathcal{G}$ and $t\geq 0$,
    $g\cdot\mathcal{S}_{t}^{W}(u_{0})\stackrel{{\scriptstyle\mathrm{law}}}{{=}}\mathcal{S}_{t}^{W}(g\cdot u_{0}).$
  2. (ii)
    (Equivariance of invariant measures.) If $\nu$ is an invariant measure for the Markov semigroup $\{\mathcal{S}_{t}^{W}\}_{t\geq 0}$ (i.e. $u_{0}\sim\nu\implies\mathcal{S}_{t}^{W}(u_{0})\sim\nu$ for all $t\geq 0$), then $g_{\#}\nu$ is also invariant for every $g\in\mathcal{G}$.
  3. (iii)
    (Equivariance of invariant fields.) If $Z$ is an invariant random field (i.e. $Z\stackrel{{\scriptstyle\mathrm{law}}}{{=}}\mathcal{S}_{t}^{W}(Z)$ for all $t\geq 0$ with $Z$ independent of $W$), then $g\cdot Z$ is also invariant for every $g\in\mathcal{G}$.
Proof.
(i) By Proposition A.2, for fixed $u_{0}$ and $g$ we have the pathwise identity
$g\cdot\mathcal{S}_{t}^{W}(u_{0})=\mathcal{S}_{t}^{W^{g}}(g\cdot u_{0}),$
where $W^{g}$ is the transformed noise. Since $W^{g}\stackrel{{\scriptstyle\mathrm{law}}}{{=}}W$, the right-hand side has the same law as $\mathcal{S}_{t}^{W}(g\cdot u_{0})$, proving (i).
(ii) Let $u_{0}\sim g_{\#}\nu$, so that $u_{0}=g\cdot\tilde{u}_{0}$ for some $\tilde{u}_{0}\sim\nu$. Then by (i),
$\mathcal{S}_{t}^{W}(u_{0})=\mathcal{S}_{t}^{W}(g\cdot\tilde{u}_{0})\ \stackrel{{\scriptstyle\mathrm{law}}}{{=}}\ g\cdot\mathcal{S}_{t}^{W}(\tilde{u}_{0}).$
Since $\nu$ is invariant, $\mathcal{S}_{t}^{W}(\tilde{u}_{0})\sim\nu$, hence $g\cdot\mathcal{S}_{t}^{W}(\tilde{u}_{0})\sim g_{\#}\nu$. Therefore $\mathcal{S}_{t}^{W}(u_{0})\sim g_{\#}\nu$ for all $t$, i.e. $g_{\#}\nu$ is invariant.
(iii) If $Z$ is invariant and independent of $W$, then $\mathcal{S}_{t}^{W}(Z)\stackrel{{\scriptstyle\mathrm{law}}}{{=}}Z$ for all $t$. Applying (i) with $u_{0}=Z$ and using independence,
$\mathcal{S}_{t}^{W}(g\cdot Z)\ \stackrel{{\scriptstyle\mathrm{law}}}{{=}}\ g\cdot\mathcal{S}_{t}^{W}(Z)\ \stackrel{{\scriptstyle\mathrm{law}}}{{=}}\ g\cdot Z,$
so $g\cdot Z$ is invariant. ∎
In particular, the (1)-invariant field corresponding to a constant harmonic function is fully invariant under the action of $\mathcal{G}$. For general $h\in\mathrm{Har}^{+}_{b}(M)$, the associated (1)-invariant fields transform equivariantly under isometries via the natural action $(g\cdot h)(x)=h(g^{-1}x),$ which corresponds to the induced action on boundary data. A similar result holds for the fluctuation Gaussian fields $G_{h}$ of Theorem 1.5 by a nearly identical calculation.